OpenStudy (anonymous):

-24>8t interval notation?

OpenStudy (anonymous):

\[ -24\gt 8t\implies -3>t\implies -\infty < t < -3 \]This is written as: \[ t\in (-\infty,-3) \]

OpenStudy (anonymous):

\[ a<x<b\implies x\in(a,b)\\ a \leq x<b\implies x\in[a,b)\\ a<x\leq b\implies x\in(a,b]\\ a \leq x \leq b\implies x\in[a,b] \]

OpenStudy (anonymous):

Notice how \(\leq\ \geq\) corresponds to \([\ ]\) and how \(\lt\ \gt\) corresponds to \((\ )\)

OpenStudy (anonymous):

When there is no lower or upper bound, you have to use \(\infty\) and \(-\infty\) to show your boundaries. They will always be used with \((\ )\) next to them because they should be excluded from the interval.